Hannibal Corp. has a cost of equity of 12%, and an after-tax cost of debt of 7%. Using market values, Hannibals debt is 40% of the value of the firm and its equity is 60% of the value of the firm. What is the weighted average cost of capital?

Dexter Corp. can borrow at 9% and is has a 23% marginal tax rate. What is the after-tax cost of debt?

Columbia Corporation has a beta of 1.6, the risk-free rate is expected to be 3.5% and the market risk premium is expected to be 5.5%. Using the Capital Asset Pricing Model, what is the after-tax cost of equity?

Quincy Corp. expects NOPLAT of $3,800,000 in the first year after the forecast period (or the first year of continuing value period. In addition, it expects growth of 2.5% during the continuing value period, and a weighted average cost of capital of 7%. Please calculate the continuing value.

St Louis Corp. expects free cash flows of $2,000,000, $3,200,000, $4,400,000, $5,800,000, $6,500,000, in years 1,2 3,4, 5, respectively. In addition, it has a continuing value of $20,000,000 at the end of year 5 and a cost of capital of 7%. Assuming year end cash flows and that these are all the cash flows for the company, what is the value of this company?
Answer #1

1. Weighted average cost of capital for Hannibal Corp is calculated using the formaula (WACC) = w(d) * k(d) + w(e) * k(e)

where w(d) denotes weight of debt

w(e) denotes weight of equity

k(d) denotes after cost of debt

k(e) denotes cost of equity

Therefore, WACC = 40% * 7% + 60% * 12% = 10%

2. After that cost of debt for Dexter Corp = cost of debt * (1 - marginal tax rate) = 9% * (1 - 23%) = 6.93%

3. Cost of equity for Columbia Corporation using WACC = risk free rate + beta* risk premium

which is equal to 3.5% + 1.6 * 5.5% = 12.3%

4. Continuing Value for Quincy Corp = NOPLAT (1) / (WACC - g) = $3800,000 / (7% - 2.5%) = $84,444,444.44

5. Value of the Company = $2,000,000 / (1+7%)^1 + $3,200,000 / (1+7%)^2 + $4,400,000 / (1+7%)^3 + $5,800,000 / (1+7%)^4 + $6,500,000 / (1+7%)^5 + $20,000,000 / (1+7%)^5 = $31,574,799.45